OpenStudy (anonymous):

what does a sideways 8 and the symbols [ and ( mean on a line chart?

OpenStudy (saifoo.khan):

[ means less than and equal to. ( means just less than.

OpenStudy (anonymous):

so i have a [ at the seven, and a straight line pointing past the ten

OpenStudy (saifoo.khan):

it means, the x is greater than and equal to seven.

OpenStudy (saifoo.khan):

can u post the photo here?

OpenStudy (anonymous):

with an arrow

OpenStudy (saifoo.khan):

yes, thts it.

OpenStudy (anonymous):

how wopuld i write that with the sideways 8?

OpenStudy (saifoo.khan):

\[x \le 7\]

OpenStudy (anonymous):

Sidways eight is \[\infty\] which means infinity.

OpenStudy (saifoo.khan):

but its better if u post a photo!

OpenStudy (saifoo.khan):

it depends how u need the answer.

OpenStudy (anonymous):

x greater than or equal to seven: \[x \in [7;\infty[ \] Or a ) after the infinity symbol. I think the convention is [ for open interval though.

OpenStudy (anonymous):

x >=7. Basically, x is anywhere from being equal to 7 to being however large. That's what the interval [7, inf) means.